What is the valuation ranges for P/E and (P/E)/g, and how did you arrive at the high and low end of your ranges?

Firm:

P/E

(P/E)/g

Lockheed Martin

18.59

28.86

Northrop Grumman

19.96

30.99

General Dynamics

15.75

24.45
Answer #1
Firm: P/E (P/E)/g
General Dynamics 15.75 24.45
Lockheed Martin 18.59 28.86
Northrop Grumman 19.96 30.99
P/E range = Highest P/E - Lowest P/E
Highest P/E = 19.96
Lowest P/E = 15.75
P/E range = 19.96 - 15.75 = 4.21
(P/E)/g range = Highest (P/E)/g - Lowest (P/E)/g
Highest (P/E)/g = 30.99
Lowest (P/E)/g = 24.45
(P/E)/g range = 30.99 - 24.45 = 6.54
We have decided the high end and low end by simply taking the ascending orders of valuation multiples. The first number is low end and last number is high end